# Bipartite entanglement extracted from multimode squeezed light generated in lossy waveguides

### v2

@@ Abstract
Usually, the internal waveguide losses for the TE- and the TM-modes are different [10]. In order to take this into account, we introduce two frequency-independent loss coefficients for the signal and idler fields $\eta_{s}$ and $\eta_{i}$, respectively. These coefficients can be parametrized as $\bar{\eta}=\frac{\eta_{s}+\eta_{i}}{2},\,r_{\eta}=\frac{\eta_{s}-\eta_{i}}{\eta_{s}+\eta_{i}}.$ In this paper, we consider three waveguides:
@@
In the limit $\bar{\eta}\rightarrow 0$, both waveguides WG1 and WG2 coincide with the lossless waveguide WG0. We numerically solve the master equations (see Appendix B) for each waveguide for different parametric gains $\Gamma$ and losses $\bar{\eta}$, $r_{\eta}$. For the resulting correlation matrices $\mathcal{D}(\Gamma,\bar{\eta},r_{\eta})$ and $\mathcal{C}(\Gamma,\bar{\eta},r_{\eta})$, we find the Mercer-Wolf, Williamson-Euler, and MSq modes. For each basis, we calculate the TMBS with the algorithms described above.
Figure 4: TMBS for the lossy waveguide WG2 (unbalanced losses) with $\bar{\eta}=5$ dB and $r=\frac{1}{3}$. (a,b,c) Covariance matrices $\sigma^{MW}$, $\sigma^{WE}$ and $\sigma^{MSq}$, respectively. (d) Absolute value and (e) phase of signal modes $\mathbf{u}_{A}$; (f) absolute value and (g) phase of idler modes $\mathbf{u}_{B}$. The blue, green, and red colors correspond to the MW-, WE-, and MSq-TMBS bases, respectively.
Fig. 2 shows results of numerical simulations for the lossless waveguide WG0, for which, as was mentioned above, there exists the optimal TMBS based on the first Schmidt mode. Fig. 2(b) demonstrates the dependence of the smallest eigenvalue $\lambda_{-}(\sigma^{S})$ (characterized by quadrature variance $\braket{(\Delta\hat{p}^{F})^{2}}$) as a function of the number of photons. The obtained TMBS is pure and corresponds to the two-mode squeezed state; therefore, entanglement and squeezing are always present in this state.
@@
In the case of unbalanced losses, choosing a wrong basis may even lead to a loss of entanglement, as it is demonstrated in Fig. 3(d) for WG2. Here, for $\bar{\eta}>2.6$ dB, the lowest eigenvalue is grater than unity, $\lambda_{-}(\sigma^{MW})>1$, which means that MW-TMBS becomes separable (see the gray region in Fig. 3(d)). For the losses that exhibit the maximum difference in the smallest eigenvalue between the three bases in Fig. 3(d), namely, $\bar{\eta}=5$ dB, we present the explicit form of the covariance matrices $\sigma^{MW}$, $\sigma^{WE}$, and $\sigma^{MSq}$, see Fig. 4(a-c), as well as the spectral and phase profiles of $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ modes, see Fig. 4(d-g). It is clearly seen that the structure of the covariance matrix and the spectral width of the modes strongly depend on the chosen basis. For example, in the MW-basis, the modes $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ have almost the same width, while in the MSq-basis their widths differ significantly. The mean number of photons, the lowest eigenvalue, as well as the negativity and purity for this special loss point are presented in Table 1 and show the importance of choosing the correct basis for a particular lossy system.
Summing up, we demonstrate how the TMBS with the maximal degree of squeezing can be created from a pulsed type-II multimode PDC process in lossy waveguides. For the obtained TMBS, we show that the maximal degree of squeezing quantifies entanglement and we introduce a basis (MSq-basis) in which the TMBS with the maximal entanglement can be constructed. We compare this optimal TMBS with the TMBS obtained via the Mercer-Wolf and the Williamson-Euler modes. By numerical simulations of lossy waveguides, we demonstrate that neither the first Mercer-Wolf mode, nor the first Williamson-Euler mode are optimal for preparing the maximally entangled state. Therefore, an intuitive experimental approach of maximizing intensity would not result in an optimal entangled TMBS.
From an experimental point of view, to find the MSq-basis, one can either implement a multiparameter optimization procedure via Eq. (7) using the two-mode squeezing $\lambda_{-}\left(\sigma(\mathbf{u}_{A},\mathbf{u}_{B})\right)$ as a target parameter or provide a diagonalization of the measured covariance matrix in order to find the proper shapes of $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ modes. Herewith, a complete classical waveguide characterization can significantly simplify the multiparameter optimization task by limiting the range of parameters.
@@ Appendix A Multimode bipartite Gaussian state
A detailed descriptions can be found in Refs. [14, 15]. Let us consider the multimode Gaussian system, consisting of two parts: the part A, which is consists of $N$ bosonic modes, denoted as $\hat{\mathbf{a}}=\big{(}\hat{a}_{0},\hat{a}_{1},\dots,\hat{a}_{N}\big{)}^{T}$ and the part B, consisting of $M$ bosonic modes $\hat{\mathbf{b}}=\big{(}\hat{b}_{0},\hat{b}_{1},\dots,\hat{b}_{M}\big{)}^{T}$. Commutation relations are bosonic $[\hat{a}_{i}(z),\hat{a}^{\dagger}_{j}(z)]=[\hat{b}_{i}(z),\hat{b}^{\dagger}_{j}(z)]=\delta_{ij}$ and $[\hat{a}_{i}(z),\hat{b}^{\dagger}_{j}(z)]=0$. The correlation matrices that fully describe the state of the full system read
@@
In addition to the creation and annihilation operators, we also need the quadrature operators. For arbitrary modes with the annihilation operator $\hat{F}$, the quadrature operators are $\hat{q}^{F}_{i}=\hat{F}^{\dagger}_{i}+\hat{F}_{i}$ and $\hat{p}^{F}_{i}=i(\hat{F}^{\dagger}_{i}-\hat{F}_{i})$ with the commutation relations $[\hat{q}^{F}_{n},\hat{p}^{F}_{m}]=2i\delta_{nm}$.
The quadrature operators allow us to introduce the covariance matrix for the full system.
@@
with the vector $\hat{\mathbf{x}}=(\hat{q}^{a}_{1},\hat{p}^{a}_{1},\dots,\hat{q}^{a}_{N},\hat{p}^{a}_{N};\hat{q}^{b}_{1},\hat{p}^{b}_{1},\dots,\hat{q}^{b}_{M},\hat{p}^{b}_{M})^{T}$, while the quadrature operators are $\hat{q}^{c}_{i}=\hat{c}^{\dagger}_{i}+\hat{c}_{i}$ and $\hat{p}^{c}_{i}=i(\hat{c}^{\dagger}_{i}-\hat{c}_{i})$ with the commutation relations $[\hat{q}^{c}_{n},\hat{p}^{d}_{m}]=2i\delta_{nm}\delta_{cd}$. Here, the upper indexes indicate arbitrary modes $c,d=(a,b)$. The elements of the covariance matrix are connected with the correlation matrices (9) as
@@ Appendix B Master equation
where the superscript $[\cdot]^{*}$ denotes the complex conjugation of a matrix. The matrix $K$ is a diagonal matrix consisting of complex wave vectors $\textrm{diag}(\kappa^{a}_{0},\dots\kappa^{a}_{N},\kappa^{b}_{0},\dots,\kappa^{b}_{N})$. Here, $\kappa^{a}_{n}=k^{a}_{n}+i\eta_{a}/2$, where $k^{a}_{n}$ is a real wave vector for mode $\hat{a}_{n}$ and $\eta_{a}$ is the loss coefficient for subsystem $a$. The $z$-dependent matrix $M(z)$ is given by
@@ Appendix C Mercer-Wolf and Williamson-Euler decompositions
are obtained via the Williamson-Euler decomposition [37, 14], namely a decomposition of the initial covariance matrix $\sigma=O_{l}\Lambda O_{r}DO_{r}^{T}\Lambda O_{l}^{T}$, where the matrices $O_{l}$ and $O_{r}$ are orthogonal and matrices $D=\text{diag}(\nu_{1},\nu_{1},\dots,\nu_{N},\nu_{N})$ and $\Lambda=\text{diag}(e^{r_{1}},e^{-r_{1}},\dots,e^{r_{N}},e^{-r_{N}})$ are diagonal; the values $r_{i}$ are sorted in descending order. Having the first vector $\mathbf{v}^{(1)}=(x^{a}_{1},y^{a}_{1},\dots,x^{a}_{N},y^{a}_{N};x^{b}_{1},y^{b}_{1},\dots,x^{b}_{M},y^{b}_{M})$ of matrix $O_{l}$, and applying the same procedure as for the MSq mode, we can build modes for each part as $\hat{A}^{WE}=\mathbf{u}^{WE}_{A}\mathbf{\hat{a}}$ and $\hat{B}^{WE}=\mathbf{u}^{WE}_{B}\mathbf{\hat{b}}$, where $\mathbf{u}^{WE}_{A}=\frac{\mathbf{v}^{a}}{|\mathbf{v}^{a}|}$ and $\mathbf{u}^{WE}_{B}=\frac{\mathbf{v}^{b}}{|\mathbf{v}^{b}|}$ with $[\mathbf{v}^{a}]_{n}=y^{a}_{n}+ix^{a}_{n}$ and $[\mathbf{v}^{b}]_{n}=y^{b}_{n}+ix^{b}_{n}$. Numerically, the Williamson-Euler decomposition can be found with the use of algorithms presented in, e.g., Refs. [35, 36, 38]. The TMBS obtained by the vectors $\mathbf{u}^{WE}_{A}$ and $\mathbf{u}^{WE}_{B}$ is further called as WE-TMBS; the corresponding covariance matrix is defined as $\sigma^{WE}\equiv\sigma(\mathbf{u}^{WE}_{A},\mathbf{u}^{WE}_{B})$.

### v3

@@ III Numerical simulations
Usually, the internal waveguide losses for the TE- and the TM-modes are different [10]. In order to take this into account, we introduce two frequency-independent loss coefficients for the signal and idler fields $\eta_{s}$ and $\eta_{i}$, respectively. We parametrize these coefficients as $\bar{\eta}=\frac{\eta_{s}+\eta_{i}}{2},\,r_{\eta}=\frac{\eta_{s}-\eta_{i}}{\eta_{s}+\eta_{i}}.$ In this paper, we consider three waveguides:
@@
In the limit $\bar{\eta}\rightarrow 0$, both waveguides WG1 and WG2 coincide with the lossless waveguide WG0. We numerically solve the master equations (see Appendix B, and Ref. 40) for each waveguide for different parametric gains $\Gamma$ and losses $\bar{\eta}$, $r_{\eta}$. For the resulting correlation matrices $\mathcal{D}(\Gamma,\bar{\eta},r_{\eta})$ and $\mathcal{C}(\Gamma,\bar{\eta},r_{\eta})$, we find the Mercer-Wolf, Williamson-Euler, and MSq modes. For each basis, we calculate the TMBS with the algorithms described above.
Figure 4: TMBS for the lossy waveguide WG2 (unbalanced losses) with $\bar{\eta}=5$ dB and $r=\frac{1}{3}$. (a,b,c) Covariance matrices $\sigma^{MW}$, $\sigma^{WE}$ and $\sigma^{MSq}$, respectively. (d) Absolute value and (e) phase of signal modes $\mathbf{u}_{A}$; (f) absolute value and (g) phase of idler modes $\mathbf{u}_{B}$. The blue, green, and red colors correspond to the MW-, WE-, and MSq-TMBS bases, respectively. Note that the phases of modes are well defined, providing the covariance matrices (a,b,c) to be in the form Eq. (6).
Fig. 2 shows results of numerical simulations for the lossless waveguide WG0, for which, as was mentioned above, there exists the optimal TMBS based on the first Schmidt mode. Note that the shape of the joint spectral intensity is similar to the results presented for the LiNbO3 based waveguides [41]. Fig. 2(b) demonstrates the dependence of the smallest eigenvalue $\lambda_{-}(\sigma^{S})$ (characterized by quadrature variance $\braket{(\Delta\hat{p}^{F})^{2}}$) as a function of the number of photons. The obtained TMBS is pure and corresponds to the two-mode squeezed state; therefore, entanglement and squeezing are always present in this state.
@@
In the case of unbalanced losses, choosing a MW-basis may even lead to a loss of entanglement, as it is demonstrated in Fig. 3(d) for WG2. Here, for $\bar{\eta}>2.6$ dB, the lowest eigenvalue is grater than unity, $\lambda_{-}(\sigma^{MW})>1$, which means that MW-TMBS becomes separable (see the gray region in Fig. 3(d)). For the losses that exhibit the maximum difference in the smallest eigenvalue between the three bases in Fig. 3(d), namely, $\bar{\eta}=5$ dB, we present the explicit form of the covariance matrices $\sigma^{MW}$, $\sigma^{WE}$, and $\sigma^{MSq}$, see Fig. 4(a-c), as well as the spectral and phase profiles of $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ modes, see Fig. 4(d-g). It is clearly seen that the covariance matrix and the spectral width of the modes strongly depend on the chosen basis. For example, in the MW-basis, the modes $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ have almost the same width, while in the MSq-basis their widths differ significantly. The mean number of photons, the lowest eigenvalue, as well as the negativity and purity for this special loss point are presented in Table 1 and show the importance of choosing the correct basis for a particular lossy system.
@@
IV Summary
Summing up, we demonstrate how to create the TMBS with the maximal degree of entanglement from a bipartite multimode Gaussian mixed state generated via pulsed type-II multimode PDC in lossy waveguides. We show that, in the obtained TMBS, the Gaussian entanglement is characterized by the degree of squeezing and show how the maximally-entangled TMBS can be constructed with the use of the MSq-basis. We compare this optimal TMBS with the TMBS obtained via the Mercer-Wolf and the Williamson-Euler modes. By numerical simulations of lossy waveguides, we demonstrate that neither the first Mercer-Wolf mode nor the first Williamson-Euler mode are optimal for preparing the maximally entangled state.
From an experimental point of view, there are three important corollaries. First, to get the maximally entangled TMBS from the type-II PDC, the measurement basis should coincide with the first mode of the MSq-basis. However, this mode does not provide the maximal number of photons, therefore, the intuitive experimental approach of maximizing intensity would not result in a maximally entangled TMBS. Second, for TMBS, we can use the two-mode squeezing as an experimental measure of entanglement. Therefore, the proper measurement basis can be found via a diagonalization of the measured covariance matrix. Alternatively, one can experimentally implement a multiparameter optimization procedure via Eq. (7) using the two-mode squeezing degree $\lambda_{-}\left(\sigma(\mathbf{u}_{A},\mathbf{u}_{B})\right)$ as a target parameter; however, this procedure requires further investigation, which is out of scope of the present paper. Third, we performed our numerical simulations only for internal PDC losses, assuming the absence of the external (transmission and detection) losses. However, as was mentioned above, frequency-dependent external losses can easily be taken into account, in particular, when the PDC light passes through spectral filters. All the procedures and arguments for building an optimal TMBS are also valid for the case of external losses and we leave its the detailed study for the future.
@@ Acknowledgements.
Data Availability Statement
Raw data were generated via the numerical solver Ref. 40.
@@ Appendix A Multimode bipartite Gaussian state
A detailed descriptions can be found in Refs. 14, 15. Let us consider the multimode Gaussian system, consisting of two parts: the part A, which is consists of $N$ bosonic modes, denoted as $\hat{\mathbf{a}}=\big(\hat{a}_{0},\hat{a}_{1},\dots,\hat{a}_{N}\big)^{T}$ and the part B, consisting of $M$ bosonic modes $\hat{\mathbf{b}}=\big(\hat{b}_{0},\hat{b}_{1},\dots,\hat{b}_{M}\big)^{T}$. Commutation relations are bosonic $\hat{a}_{i}(z),\hat{a}^{\dagger}_{j}(z)]=[\hat{b}_{i}(z),\hat{b}^{\dagger}_{j}(z)]=\delta_{ij}$ and $\hat{a}_{i}(z),\hat{b}^{\dagger}_{j}(z)]=0$. The correlation matrices that fully describe the state of the full system read
@@
Alternatevely, the state of the full system can be described by the covariance matrix $\Sigma$ with matrix elements
@@
where the vector $\hat{\mathbf{x}}=(\hat{q}^{a}_{1},\hat{p}^{a}_{1},\dots,\hat{q}^{a}_{N},\hat{p}^{a}_{N};\hat{q}^{b}_{1},\hat{p}^{b}_{1},\dots,\hat{q}^{b}_{M},\hat{p}^{b}_{M})^{T}$, while the quadrature operators are $\hat{q}^{c}_{i}=\hat{c}^{\dagger}_{i}+\hat{c}_{i}$ and $\hat{p}^{c}_{i}=i(\hat{c}^{\dagger}_{i}-\hat{c}_{i})$ with the commutation relations $\hat{q}^{c}_{n},\hat{p}^{d}_{m}]=2i\delta_{nm}\delta_{cd}$. Here, the upper indexes indicate arbitrary modes $c,d=(a,b)$. The elements of the covariance matrix are connected with the correlation matrices (11) as
@@
In addition to the creation and annihilation operators, we also need the quadrature operators for new broadband modes. For an arbitrary mode with the annihilation operator $\hat{F}_{n}$, the quadrature operators are $\hat{q}^{F}_{n}=\hat{F}^{\dagger}_{n}+\hat{F}_{n}$ and $\hat{p}^{F}_{n}=i(\hat{F}^{\dagger}_{n}-\hat{F}_{n})$ with the commutation relations $\hat{q}^{F}_{n},\hat{p}^{F}_{m}]=2i\delta_{nm}$.
@@ Appendix B Master equation
where the superscript $\cdot$$^{*}$ denotes the complex conjugation of a matrix. The matrix $K$ is a diagonal matrix consisting of complex wave vectors $\textrm{diag}(\kappa^{a}_{0},\dots\kappa^{a}_{N},\kappa^{b}_{0},\dots,\kappa^{b}_{N})$. Here, $\kappa^{a}_{n}=k^{a}_{n}+i\eta_{a}/2$, where $k^{a}_{n}$ is a real wave vector for mode $\hat{a}_{n}$ and $\eta_{a}$ is the loss coefficient for subsystem $a$. The $z$-dependent matrix $M(z)$ is given by
@@ Appendix C Mercer-Wolf and Williamson-Euler decompositions
are obtained via the Williamson-Euler decomposition [37, 14], namely a decomposition of the initial covariance matrix $\sigma=O_{l}\Lambda O_{r}DO_{r}^{T}\Lambda O_{l}^{T}$, where the matrices $O_{l}$ and $O_{r}$ are orthogonal and matrices $D=\text{diag}(\nu_{1},\nu_{1},\dots,\nu_{N},\nu_{N})$ and $\Lambda=\text{diag}(e^{r_{1}},e^{-r_{1}},\dots,e^{r_{N}},e^{-r_{N}})$ are diagonal; the values $r_{i}$ are sorted in descending order. Having the first vector $\mathbf{v}^{(1)}=(x^{a}_{1},y^{a}_{1},\dots,x^{a}_{N},y^{a}_{N};x^{b}_{1},y^{b}_{1},\dots,x^{b}_{M},y^{b}_{M})$ of matrix $O_{l}$, and applying the same procedure as for the MSq mode, we can build modes for each part as $\hat{A}^{WE}=\mathbf{u}^{WE}_{A}\mathbf{\hat{a}}$ and $\hat{B}^{WE}=\mathbf{u}^{WE}_{B}\mathbf{\hat{b}}$, where $\mathbf{u}^{WE}_{A}=\frac{\mathbf{v}^{a}}{|\mathbf{v}^{a}|}$ and $\mathbf{u}^{WE}_{B}=\frac{\mathbf{v}^{b}}{|\mathbf{v}^{b}|}$ with $\mathbf{v}^{a}]_{n}=y^{a}_{n}+ix^{a}_{n}$ and $\mathbf{v}^{b}]_{n}=y^{b}_{n}+ix^{b}_{n}$. Numerically, the Williamson-Euler decomposition can be found with the use of algorithms presented in, e.g., Refs. [35, 36, 38]. The TMBS obtained by the vectors $\mathbf{u}^{WE}_{A}$ and $\mathbf{u}^{WE}_{B}$ is further called as WE-TMBS; the corresponding covariance matrix is defined as $\sigma^{WE}\equiv\sigma(\mathbf{u}^{WE}_{A},\mathbf{u}^{WE}_{B})$.